what is the length of the hypotenuse of a right triangle that has legs with lengths of 5 and 4?
Answer by nerdybill(7384)

what is the length of the hypotenuse of a right triangle that has legs with lengths of 5 and 4?
apply the Pythagorean Theorem:
Let h = hypotenuse
then
h^2 = 5^2 + 4^2
h^2 = 25 + 16
h^2 = 41
h = sqrt(41)
h = 6.40